OpenStudy (anonymous):

Add. Assume that all variables represent positive real numbers. 6sqrt(3x^3) + 3xsqrt(12x)

OpenStudy (anonymous):

\[=6x \sqrt{3x}+6x \sqrt{3x}=12x \sqrt{3x}\]